OpenStudy (melodycrazysouthern):

What is the fraction for the following decimal expansion? 0.125

OpenStudy (sapphiremoon):

125/1000 (divide numerator and denominator by 5) = 25/200 (divide by 5 again) = 5/40 (divide by 5 again) = 1/8. (You could also just divide by 125 in the first place, but this shows how we get there better, IMO).

OpenStudy (melodycrazysouthern):

ooooohhhhhh that makes alot of sense